In newton's ring experiment the diameter of nineth bright ring changes from 0.3cm to 0.26m.when a liquid is introduced between the lens and the glass plate.what will be the refractive index of the liquid

1
Expert's answer
2022-01-11T14:37:51-0500

The refractive index of the liquid is


"n=\dfrac{D^2_\text{air}}{D^2_\text{liquid}}=\dfrac{0.3^2}{0.26^2}=1.33."